how much of a 7% salt solution and a 1% salt solution is needed to make 1 liter of 3% salt solution?

Answer by checkley79(3341)

.07X+.01(3-X)=.03*3
.07X+.03-.01X=.09
.06X=.09-.03
.06X=.06
X=.06/.06
X=1 LITER OF 7% SOLUTION IS USED.
3-1=2 LITERS OF 1% SOLUTION IS USED.
PROOF:
.07*1+.01*2=.03*3
.07+.02=.09
.09=.09